sum of a number and its double is 9.Find the number.
Answer by lenny460(1073)

Sum of a number and its double is 9. Find the number ?
Let the number be x.
Its double is 2x.
Therefore:
x + 2x = 9
3x = 9
Divide each side by 3
3x/3 = 9/3
x = 3



Answer:
3









Lennox Obuong
Algebra Tutor
Nairobi, Kenya
Email: obuong3@aol.com